# FaithLM: Towards Faithful Explanations for Large Language Models

### v3

@@ Appendix G Details of Prompts Usage in FaithLM
Contrary Explanations Generation Please generate one example of obtaining the opposite meaning from a given sentence. Make sure you output sentences only. Sentences:{derived explanation}.
Table 8: The example of prompts that are given to two explainer LLMs and LLM agent for contrary explanation.
Appendix H Trajectory System Prompts Usage in FaithLM
@@
H.1 Trigger-oriented Trajectory Prompt
@@
H.2 NL Explanation-oriented Trajectory Prompt
@@
Appendix I Case Studies of FaithLM
I.1 Explanations Generation
@@
I.2 Explanation Trigger Prompts
@@
I.3 Contrary Explanation
We demonstrate results on contrary explanations leveraged in Fidelity Evaluator. The results in Figure 17 show that the contrary explanation obtains the opposite meaning of the given explanation.
Case Study 1 Question: The child wanted to test electricity, so what did he feel as a result? Choices: [choice] cause fire [choice] shock [choice] short fuse [choice] kill if used improperly [choice] produce heat Ground Truth: shock LLM Answer: cause fire (wrong) Given Explanation: The model likely incorrectly associated testing electricity with igniting a fire, rather than understanding that direct contact can cause an electric shock. It failed to comprehend the potential outcomes of unsafe electrical contact. Contrary Explanation: The model correctly predicted that testing electricity would not cause a fire. It likely fully understood that direct contact with electricity can cause a shock rather than ignite a fire. Case Study 2 Question: Where would you put coins if you want to bring them with you? Choices: [choice] medium [choice] water fountain [choice] wishing well [choice] desk [choice] purse Ground Truth: purse LLM Answer: purse (correct) Given Explanation: A purse is a personal item that people often carry with them when going places. It has compartments to store small items like coins, so putting coins in your purse allows you to easily bring them along wherever you go. Contrary Explanation: The purse is not a good place to put coins if you don’t want to bring them with you, because purses are designed for other items, not coins.
Figure 17: Case Study: Contrary Explanation (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)
I.4 Limitations of Learning FaithLM
One significant limitation of FaithLM associated with the carbon emissions during the experiments. The extensive computational power required for optimizing FaithLM leads to considerable energy consumption, which, in turn, results in a significant carbon footprint. As the demand for more sophisticated LLMs continues to grow, so does their environmental impact. This limitation underscores the urgent need to explore and adopt more sustainable practices and technologies in the development and learning FaithLM with fewer optimization steps to mitigate their ecological footprint.

### v4

@@ Appendix I Details of Prompts Usage in FaithLM
contrary hints Generation Please generate one example of obtaining the opposite meaning from a given sentence. Make sure you output sentences only. Sentences:{derived explanation}.
Table 8: The example of prompts that are given to two explainer LLMs and LLM agent for contrary hint.
Appendix J Trajectory System Prompts Usage in FaithLM
@@
J.1 Trigger-oriented Trajectory Prompt
@@
J.2 NL Explanation-oriented Trajectory Prompt
@@
Appendix K Case Studies of FaithLM
K.1 Explanations Generation
@@
K.2 Explanation Trigger Prompts
@@
K.3 contrary hint
We demonstrate results on contrary hints leveraged in Fidelity Evaluator. The results in Figure 18 show that the contrary hint obtains the opposite meaning of the given explanation.
Case Study 1 Question: The child wanted to test electricity, so what did he feel as a result? Choices: [choice] cause fire [choice] shock [choice] short fuse [choice] kill if used improperly [choice] produce heat Ground Truth: shock LLM Answer: cause fire (wrong) Given Explanation: The model likely incorrectly associated testing electricity with igniting a fire, rather than understanding that direct contact can cause an electric shock. It failed to comprehend the potential outcomes of unsafe electrical contact. contrary hint: The model correctly predicted that testing electricity would not cause a fire. It likely fully understood that direct contact with electricity can cause a shock rather than ignite a fire. Case Study 2 Question: Where would you put coins if you want to bring them with you? Choices: [choice] medium [choice] water fountain [choice] wishing well [choice] desk [choice] purse Ground Truth: purse LLM Answer: purse (correct) Given Explanation: A purse is a personal item that people often carry with them when going places. It has compartments to store small items like coins, so putting coins in your purse allows you to easily bring them along wherever you go. contrary hint: The purse is not a good place to put coins if you don’t want to bring them with you, because purses are designed for other items, not coins.
Figure 18: Case Study: contrary hint (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)